OpenStudy (anonymous):

Which of the following is the equation of a line perpendicular to the line y=-3/2x+4 , passing through the point (3,9)? (will post answer choices in a sec!!) **Please explain!! thanks y'all :)

OpenStudy (anonymous):

answer choices A,B,C,D (top to bottom)

jimthompson5910 (jim_thompson5910):

the slope of y=-3/2x+4 is ???

OpenStudy (anonymous):

-3/2?

jimthompson5910 (jim_thompson5910):

Flip the fraction to get -2/3 then flip the sign to get +2/3 or just 2/3

jimthompson5910 (jim_thompson5910):

so the perpendicular slope is 2/3

OpenStudy (anonymous):

cuz perpendicular lines have opposite reciprocal slopes right?? so the slope's 2/3 :) did i get that right? the concept i mean?

jimthompson5910 (jim_thompson5910):

you are absolutely correct

OpenStudy (anonymous):

so i need to solve each equation (in the answer choices) to see which has a slope of 2/3 ??

jimthompson5910 (jim_thompson5910):

y = mx+b y = 2/3x + b ... Plug in m = 2/3 9 = (2/3)(3) + b ... Plug in (3,9) 9 = 6/3 + b ... Multiply 9 = 2 + b Keep going to solve for b

OpenStudy (anonymous):

so A is out cuz i did this.